3. If BEJO is a square, mβˆ π΅π‘‚π½ = 4π‘₯ βˆ’ 6 and 𝐡𝑂= 2x- 8, find 𝐽𝑂.




1
Expert's answer
2021-04-29T16:56:46-0400

if BEJO is a square, angle BOJ =

"4x-6"

then

"4x-6=90"

"4x=96"

"x=24"

"BO=2*48-8\n=40" units